OpenStudy (chycora):

Irma needs to rent a car for a business trip from Phoenix to San Jose (720 miles one way). She will drive round trip and will be gone a total of 8 days. The car rental company charges $65.00 per day and allows for 1,600 miles for Irmas's trip. There will be an additional charge of $0.24 per mile over the 1,600 miles allowed. How many total miles can she drive without exceeding her budget of $600.00? A. 1,773 miles B. 333 miles C. 1,053 miles D. 1,933 miles

OpenStudy (phi):

First, if it costs $65 per day, how much does 8 days cost ?

OpenStudy (phi):

you could write down 65 for day 1 + 65 for day 2 + and so on but most people would use multiplication

OpenStudy (anonymous):

Daily + First 1600 + Excess = 600 $8 * 65 + $0 * 1600 + $0.24 * E = $600 $520 + $0.24 * E = $600 $0.24 * E = $80 E = 80/0.24 = 333 and change

OpenStudy (chycora):